# A Reinforcement Learning Method to Factual and Counterfactual Explanations for Session-based Recommendation

### v1

@@ 1. Introduction
Session-based recommendation (SR) has become integral in industries such as e-commerce and multimedia streaming, where understanding user intent within a single interaction session is crucial (Wang et al., 2021). Most existing research on SR aims to enhance recommendation accuracy by employing complex methods like Markov Chains (Stamenkovic et al., 2022) and deep learning methods like recurrent neural networks (RNN) (Hidasi et al., 2015; Quadrana et al., 2017), attention mechanisms (Liu et al., 2018), and graph neural networks (GNN) (Wu et al., 2019). However, these deep learning approaches often lack transparency, making it difficult for users to understand the reasoning behind the recommendations. This opacity can ultimately reduce user experience and satisfaction. Therefore, it is essential to develop SR models that not only prioritize accuracy but also offer meaningful explanations for their recommendations.
Typically, effective explanations should fulfill two key purposes: building trust by enabling a normative evaluation of algorithmic decisions, and providing actionable recourse for users to improve future outcomes (Selbst and Barocas, 2018; Galhotra et al., 2021). To achieve these objectives, existing methods in explainable artificial intelligence (XAI) can be broadly categorized along two dimensions: the timing of explanation generation, spanning post-hoc explainable methods and model-intrinsic methods (Mittelstadt et al., 2019), and the dependency on model internals, ranging from model-dependent to model-agnostic approaches (Guidotti et al., 2018).
@@
In sequential data scenarios, recent efforts have focused on improving explainability through various approaches. These include the integrated path and embedding method (Hou and Shi, 2019) and the knowledge graph-based approach RSL-GRU (Cui et al., 2022), both of which are post-hoc methods, as well as the model-intrinsic path-based framework REKS (Wu et al., 2023). However, these existing methods face two significant limitations: they often require additional data resources such as knowledge graphs and fail to provide true explanations. A true explanation, as defined by (Xu et al., 2021; Peake and Wang, 2018), identifies the specific components of past inputs that significantly influence the recommendation outcome, thereby clarifying which items drive the system’s decision. By capturing the causal relationships between inputs and outputs, true explanations help avoid generating incorrect or misleading interpretations. Therefore, in this work, we focus on generating true explanations for SR through a post-hoc, model-agnostic framework.
@@
Counterfactual-thinking based explanations have been widely applied in recommender systems (RSs) (Tan et al., 2021; Gao et al., 2024). These approaches aim to provide true explanations by identifying minimal perturbations that can alter the recommended outputs. However, when applied to sequential data, current methods are subject to two notable challenges. First, most existing explanations primarily target non-sequential data. Second, while some progress has been made for sequential data—such as (Xu et al., 2021), which explored true explanations for sequential recommendations by identifying influential items through causal dependencies using counterfactual data, there are still several weaknesses: (1) It lacks practical insights that could be translated into actionable interventions (i.e., modifying user interactions or preferences to influence recommendation outcomes). (2) The variational autoencoder (VAE) they employed for perturbation shows limitations in capturing sequential patterns effectively. (3) Their approximate methods for item substitution in recommendation lists may lead to potential biases.
@@
While counterfactual reasoning identifies only the most critical information where removing key sub-items would alter recommendation predictions, factual reasoning retains sufficient sub-items to maintain the same prediction, though potentially including redundancies. Based on the definitions of necessity and sufficiency conditions ($X$ is necessary for $Y$ if and only if $Y\rightarrow X$ and $X$ is sufficient for $Y$ if and only if $X\rightarrow Y$), counterfactual reasoning corresponds to necessary conditions, as $Y\rightarrow X$ is equivalent to $\neg X\rightarrow\neg Y$ (removing sub-items changes the result). In contrast, factual reasoning represents sufficient conditions, as retaining sub-items preserves the result. Combining both approaches can balance necessity and sufficiency, as purely sufficient explanations may include redundancies while purely necessary explanations might miss important items. This balance has been frequently explored in explaining machine learning models (Galhotra et al., 2021) and GNN-based recommendation models (Tan et al., 2022; Chen et al., 2022b), yet it remains unexplored in the context of SR scenarios.
@@
In this context, we first formally define factual and counterfactual reasoning in SR. As illustrated in Figure 1, intuitively, the sufficient (i.e., factual) conditions for an item to be recommended to a user are the set of items that enable that item to be included in the final list of top-$K$ recommended items. A factual explanation might be: “You were recommended item $Y$ because you interacted with item $X$”. In contrast, the necessary (i.e., counterfactual) conditions are the set of items without which that item would not appear in the final list of top-$K$ recommended items. A counterfactual explanation example could be: “If you had not interacted with item $X$, you would not have been recommended item $Y$”. Generating sufficient and necessary model-agnostic explanations for existing SR models is valuable but also faces two major challenges: (1) Effectively modeling sequential data to capture dependencies between items, rather than treating them as isolated entities. (2) Balancing factual and counterfactual objectives (i.e., sufficiency and necessity) in both modeling and evaluation ensures that explanations account for both perspectives effectively.
@@ 2.1. Explainable Recommendation
Explainable Recommendation aims to enhance user satisfaction and system transparency (Zhang et al., 2020), typically classified into model-dependent and model-agnostic methods. The former (Shulman and Wolf, 2020; Lin et al., 2000; Vig et al., 2009) integrates explanations directly into the recommendation process but often increases complexity and may reduce overall explainability. In contrast, the latter (Verma et al., 2020; Peake and Wang, 2018) generates explanations independently, offering greater flexibility. Some studies have explored explainability for sequential data (Hou and Shi, 2019; Cui et al., 2022; Xu et al., 2021), but mainly focused on sequential recommendation with explicit user information, rather than SR where users are anonymous. Additionally, many of them require supplementary data, such as knowledge graph. In SR, REKS (Wu et al., 2023) introduces a path-based generic explainable framework, though it still requires extra knowledge graph information. To summarize, while few studies have explored explainable SR, most of them for sequential data depend on side information, and fail to provide true explanations (identifying causally influential inputs) and actionable explanations (enabling users to adjust interactions to change outcomes), limiting their effectiveness in enhancing transparency.
@@ 2.2. Factual and Counterfactual Explanation
Factual explanations, like LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017), highlight key input components that sustain a model’s prediction by removing less important information. In recommendation, especially those utilizing GNNs (Chen et al., 2022b), they enhance interpretability by identifying crucial interactions that influence recommendations. Conversely, counterfactual reasoning focuses on how changes in input can alter model predictions. It has been applied to fairness (Ge et al., 2022; Kusner et al., 2017), augment training samples (Zhang et al., 2021; Chen et al., 2022c), causality learning (Song et al., 2023) and casual explanation (Xu et al., 2021). While factual and counterfactual reasoning are widely applied, there is a need to combine their strengths to identify sufficient and necessary conditions to explain SR with sequential properties, and to use these explanations to improve recommendation accuracy.
@@ 2.3. Contrastive Learning
Recent advancements in recommendation have utilized contrastive learning (CL) to enhance model performance and robustness. In non-sequential recommendations, CL leverages static features, improving data representations without relying on temporal dynamics (Zhou et al., 2023; Yu et al., 2022). In sequential recommendations, studies such as (Zhang et al., 2024; Yang et al., 2023; Xie et al., 2022; Chen et al., 2022a; Liu et al., 2021) demonstrate the application of CL to sequential data, where multiple temporal views of user interactions are contrasted to refine the prediction models. In our study, we, for the first time, exploit factual and counterfactual explanations to provide high-quality positive and negative samples, which enhance the accuracy of SR models by taking the advantages of contrastive learning.
@@ 3.1. Task Formulation
For each session $\mathcal{S}$, we use a SR model $f_{R}$ (e.g., NARM (Li et al., 2017)) to generate embeddings for the entire session and individual items within it. $f_{R}$ produces a top-$K$ recommendation list for the next-item prediction task:
@@
(2) $\footnotesize\begin{array}[]{ll}\operatorname{minimize}&C(M)\\ \text{s.t.,}&i^{*}\in\mathbb{I}_{\mathcal{S}\odot M}^{K}\\ &i^{*}\notin\mathbb{I}_{\mathcal{S}\backslash(\mathcal{S}\odot M)}^{K}\\ &M_{i}\in\{0,1\},\quad\forall i=1,2,\ldots,|\mathcal{S}|.\end{array}$
@@ 3.2. Analysis of the Min-Explanation Problem
Time Complexity Analysis. For a session $\mathcal{S}$ with $|\mathcal{S}|$ items, the solution space comprises all possible subsets of these items, resulting in $2^{|\mathcal{S}|}$ combinations. The number of potential solutions grows exponentially with $|\mathcal{S}|$, making any algorithm that exhaustively searches for the optimal solution incur exponential time complexity. This renders the problem computationally infeasible for large $|\mathcal{S}|$. We demonstrate the computational hardness of the Min-Explanation problem by reducing it from the NP-Complete MIN-FEATURES problem (Davies and Russell, 1994).
@@
Using Reinforcement Learning to Solve the Min-Explanation Problem. In recent years, novel approaches leveraging deep reinforcement learning (RL) have emerged as effective tools for addressing combinatorial optimization problems (Mazyavkina et al., 2021). These methods have shown remarkable success in solving classical problems such as the Traveling Salesman Problem (TSP) (Bello et al., 2017). Building upon these advances, we propose a RL-based solution for the Min-Explanation problem within a Markov Decision Process (MDP) framework. Our choice is motivated by three key advantages.(1) Selecting the optimal subset of items for explanation is inherently a combinatorial optimization problem. Traditional methods struggle with the exponential search space, especially for longer sessions, making traditional exhaustive or heuristic methods computationally infeasible. RL, on the other hand, can effectively navigate this space by learning a policy for sequential decision-making. (2) SR models deal with user interaction sequences. RL is particularly well-suited to capture the temporal dynamics and dependencies within a session. (3) The Min-Explanation problem involves balancing multiple objectives: minimizing explanation complexity while satisfying both factual and counterfactual conditions. RL enables the design of a reward function that effectively balances these competing objectives. Beyond generating explanations, we also aim to use these explanations to further enhance the accuracy of SR model $f_{R}$.
@@ 4.1. Explanation Generation Module
Agent. The policy network, modeled as a MLP, integrates both session and explanation information. On this basis, a sigmoid function is applied to compute the probability of selecting the next item, following the approach suggested by recent literature (Wu et al., 2023). The policy $\pi(a_{t}\mid s_{t})$ guiding action selection is defined as:
@@
where $\sigma(\cdot)$ is the sigmoid function,$\mathbf{W}$ is a trainable parameter matrix, and $A_{t}$ is a vector representing the probability of each action. As this is a binary classification problem, the action $a_{t}$ is determined based on the output probability:
@@
1 _$\_\rightarrow v_{2}\rightarrow v_{3}\rightarrow v_{4}\rightarrow v_{5}$setminus $1$
2 $v_{1}\rightarrow\_\rightarrow v_{3}\rightarrow v_{4}\rightarrow v_{5}$ $v_{1}$ 00
3 $v_{1}\rightarrow v_{2}\rightarrow\_\rightarrow v_{4}\rightarrow v_{5}$ $v_{1}$ 00
4 $v_{1}\rightarrow v_{2}\rightarrow v_{3}\rightarrow\_\rightarrow v_{5}$ $v_{1}$ $1$
5 $v_{1}\rightarrow v_{2}\rightarrow v_{3}\rightarrow v_{4}\rightarrow\_$ $v_{1}\rightarrow v_{4}$ $1$
Final - $v_{1}\rightarrow v_{4}\rightarrow v_{5}$
A Toy Example of the MDP process. To illustrate the execution process of our FCESR framework, consider a target session “$v_{1}\rightarrow v_{2}\rightarrow v_{3}\rightarrow v_{4}\rightarrow v_{5}$”. The task involves traversing the session sequentially. Starting with an empty CF2 set $\mathcal{S}^{*}$, the agent decides at each step whether to include the current item in $\mathcal{S}^{*}$ based on the action $a_{t}$. At step 1, the agent begins with $v_{1}$ and chooses to include it in the explanation since $a_{1}=1$, updating $\mathcal{S}^{*}$ to $\{v_{1}\}$. At step 2, the agent moves to $v_{2}$ but decides not to include it since $a_{2}=0$, leaving $\mathcal{S}^{*}$ unchanged. In step 3, the agent processes $v_{3}$ and again decides not to include it in the explanation since $a_{3}=0$. At step 4, the agent evaluates $v_{4}$ and chooses to add it to $S^{*}$ since $a_{4}=1$, updating the explanation to $\{v_{1},v_{4}\}$. Finally, at step 5, the agent considers $v_{5}$ and includes it in the explanation since $a_{5}=1$, resulting in the final CF2 set $\mathcal{S}^{*}=\{v_{1},v_{4},v_{5}\}$. The action selection at each step is determined by a policy network, which integrates the embeddings of the current CF2. This example demonstrates how our framework effectively reduces the original session to its key components through a sequential decision-making process.

### v2

@@ 1. Introduction
Session-based recommendation (SR) has become integral in industries such as e-commerce and multimedia streaming, where understanding user intent within a single interaction session is crucial (wang2021survey). Most existing research on SR aims to enhance recommendation accuracy by employing complex methods like Markov Chains (Stamen2022) and deep learning methods like recurrent neural networks (RNN) (hidasi2015session; quad17HRNN), attention mechanisms (Liu2018STAMP), and graph neural networks (GNN) (wu2019session). However, these deep learning approaches often lack transparency, making it difficult for users to understand the reasoning behind the recommendations. This opacity can ultimately reduce user experience and satisfaction. Therefore, it is essential to develop SR models that not only prioritize accuracy but also offer meaningful explanations for their recommendations.
Typically, effective explanations should fulfill two key purposes: building trust by enabling a normative evaluation of algorithmic decisions, and providing actionable recourse for users to improve future outcomes (selbst2018intuitive; 2021generalexp). To achieve these objectives, existing methods in explainable artificial intelligence (XAI) can be broadly categorized along two dimensions: the timing of explanation generation, spanning post-hoc explainable methods and model-intrinsic methods (mittelstadt2019explaining), and the dependency on model internals, ranging from model-dependent to model-agnostic approaches (2018expsurvey).
@@
In sequential data scenarios, recent efforts have focused on improving explainability through various approaches. These include the integrated path and embedding method (2019houESRKG) and the knowledge graph-based approach RSL-GRU (Cui2022RSL), both of which are post-hoc methods, as well as the model-intrinsic path-based framework REKS (wu2023generic). However, these existing methods face two significant limitations: they often require additional data resources such as knowledge graphs and fail to provide true explanations. A true explanation, as defined by (xu2021learning; peake2018kdd), identifies the specific components of past inputs that significantly influence the recommendation outcome, thereby clarifying which items drive the system’s decision. By capturing the causal relationships between inputs and outputs, true explanations help avoid generating incorrect or misleading interpretations. Therefore, in this work, we focus on generating true explanations for SR through a post-hoc, model-agnostic framework.
@@
Counterfactual-thinking based explanations have been widely applied in recommender systems (RSs) (tan2021counterfactual; gao2024causal). These approaches aim to provide true explanations by identifying minimal perturbations that can alter the recommended outputs. However, when applied to sequential data, current methods are subject to two notable challenges. First, most existing explanations primarily target non-sequential data. Second, while some progress has been made for sequential data—such as (xu2021learning), which explored true explanations for sequential recommendations by identifying influential items through causal dependencies using counterfactual data, there are still several weaknesses: (1) It lacks practical insights that could be translated into actionable interventions (i.e., modifying user interactions or preferences to influence recommendation outcomes). (2) The variational autoencoder (VAE) they employed for perturbation shows limitations in capturing sequential patterns effectively. (3) Their approximate methods for item substitution in recommendation lists may lead to potential biases.
@@
While counterfactual reasoning identifies only the most critical information where removing key sub-items would alter recommendation predictions, factual reasoning retains sufficient sub-items to maintain the same prediction, though potentially including redundancies. Based on the definitions of necessity and sufficiency conditions ($X$ is necessary for $Y$ if and only if $Y\rightarrow X$ and $X$ is sufficient for $Y$ if and only if $X\rightarrow Y$), counterfactual reasoning corresponds to necessary conditions, as $Y\rightarrow X$ is equivalent to $\neg X\rightarrow\neg Y$ (removing sub-items changes the result). In contrast, factual reasoning represents sufficient conditions, as retaining sub-items preserves the result. Combining both approaches can balance necessity and sufficiency, as purely sufficient explanations may include redundancies while purely necessary explanations might miss important items. This balance has been frequently explored in explaining machine learning models (2021generalexp) and GNN-based recommendation models (tan2022learning; chen2022grease), yet it remains unexplored in the context of SR scenarios.
@@
In this context, we first formally define factual and counterfactual reasoning in SR. As illustrated in Figure 1, intuitively, the sufficient (i.e., factual) conditions for an item to be recommended to a user are the set of items that enable that item to be included in the final list of top-$K$ recommended items. A factual explanation might be: “You were recommended item $Y$ because you interacted with item XX”. In contrast, the necessary (i.e., counterfactual) conditions are the set of items without which that item would not appear in the final list of top-$K$ recommended items. A counterfactual explanation example could be: “If you had not interacted with item $X$, you would not have been recommended item YY”. Generating sufficient and necessary model-agnostic explanations for existing SR models is valuable but also faces two major challenges: (1) Effectively modeling sequential data to capture dependencies between items, rather than treating them as isolated entities. (2) Balancing factual and counterfactual objectives (i.e., sufficiency and necessity) in both modeling and evaluation ensures that explanations account for both perspectives effectively.
@@ 2.1. Explainable Recommendation
Explainable Recommendation aims to enhance user satisfaction and system transparency (zhang2020explainable), typically classified into model-dependent and model-agnostic methods. The former (shulman2020meta; lin2000collaborative; vig2009) integrates explanations directly into the recommendation process but often increases complexity and may reduce overall explainability. In contrast, the latter (verma2020counterfactual; peake2018kdd) generates explanations independently, offering greater flexibility. Some studies have explored explainability for sequential data (2019houESRKG; Cui2022RSL; xu2021learning), but mainly focused on sequential recommendation with explicit user information, rather than SR where users are anonymous. Additionally, many of them require supplementary data, such as knowledge graph. In SR, REKS (wu2023generic) introduces a path-based generic explainable framework, though it still requires extra knowledge graph information. To summarize, while few studies have explored explainable SR, most of them for sequential data depend on side information, and fail to provide true explanations (identifying causally influential inputs) and actionable explanations (enabling users to adjust interactions to change outcomes), limiting their effectiveness in enhancing transparency.
@@ 2.2. Factual and Counterfactual Explanation
Factual explanations, like LIME (ribeiro2016should) and SHAP (lundberg2017unified), highlight key input components that sustain a model’s prediction by removing less important information. In recommendation, especially those utilizing GNNs (chen2022grease), they enhance interpretability by identifying crucial interactions that influence recommendations. Conversely, counterfactual reasoning focuses on how changes in input can alter model predictions. It has been applied to fairness (ge2022explainable; kusner2017counterfactual), augment training samples (zhang2021causerec; chen2022data), causality learning (song2023counterfactual) and casual explanation (xu2021learning). While factual and counterfactual reasoning are widely applied, there is a need to combine their strengths to identify sufficient and necessary conditions to explain SR with sequential properties, and to use these explanations to improve recommendation accuracy.
@@ 2.3. Contrastive Learning
Recent advancements in recommendation have utilized contrastive learning (CL) to enhance model performance and robustness. In non-sequential recommendations, CL leverages static features, improving data representations without relying on temporal dynamics (zhou2023contrastive; Yu2022SelfSupervisedLF). In sequential recommendations, studies such as (zhang2024tois; Yang2023TheWebConf; Xie2022ContrastiveLF; Chen2022IntentCL; Liu2021ContrastiveSS) demonstrate the application of CL to sequential data, where multiple temporal views of user interactions are contrasted to refine the prediction models. In our study, we, for the first time, exploit factual and counterfactual explanations to provide high-quality positive and negative samples, which enhance the accuracy of SR models by taking the advantages of contrastive learning.
@@ 3.1. Task Formulation
For each session $\mathcal{S}$, we use a SR model $f_{R}$ (e.g., NARM (li2017neural)) to generate embeddings for the entire session and individual items within it. $f_{R}$ produces a top-$K$ recommendation list for the next-item prediction task:
@@
(2) $\footnotesize\begin{array}[]{ll}\operatorname{minimize}&C(M)\\ \text{s.t.,}&i^{*}\in\mathbb{I}_{\mathcal{S}\odot M}^{K}\\ &i^{*}\notin\mathbb{I}_{\mathcal{S}\backslash(\mathcal{S}\odot M)}^{K}\\ &M_{i}\in\{0,1\},\qquad\forall i=1,2,\ldots,|\mathcal{S}|.\end{array}$
@@ 3.2. Analysis of the Min-Explanation Problem
Time Complexity Analysis. For a session $\mathcal{S}$ with $|\mathcal{S}|$ items, the solution space comprises all possible subsets of these items, resulting in $2^{|\mathcal{S}|}$ combinations. The number of potential solutions grows exponentially with $|\mathcal{S}|$, making any algorithm that exhaustively searches for the optimal solution incur exponential time complexity. This renders the problem computationally infeasible for large $|\mathcal{S}|$. We demonstrate the computational hardness of the Min-Explanation problem by reducing it from the NP-Complete MIN-FEATURES problem (davies1994np).
@@
Using Reinforcement Learning to Solve the Min-Explanation Problem. In recent years, novel approaches leveraging deep reinforcement learning (RL) have emerged as effective tools for addressing combinatorial optimization problems (mazyavkina2021reinforcement). These methods have shown remarkable success in solving classical problems such as the Traveling Salesman Problem (TSP) (NCO2017). Building upon these advances, we propose a RL-based solution for the Min-Explanation problem within a Markov Decision Process (MDP) framework. Our choice is motivated by three key advantages.(1) Selecting the optimal subset of items for explanation is inherently a combinatorial optimization problem. Traditional methods struggle with the exponential search space, especially for longer sessions, making traditional exhaustive or heuristic methods computationally infeasible. RL, on the other hand, can effectively navigate this space by learning a policy for sequential decision-making. (2) SR models deal with user interaction sequences. RL is particularly well-suited to capture the temporal dynamics and dependencies within a session. (3) The Min-Explanation problem involves balancing multiple objectives: minimizing explanation complexity while satisfying both factual and counterfactual conditions. RL enables the design of a reward function that effectively balances these competing objectives. Beyond generating explanations, we also aim to use these explanations to further enhance the accuracy of SR model $f_{R}$.
@@ 4.1. Explanation Generation Module
Agent. The policy network, modeled as a MLP, integrates both session and explanation information. On this basis, a sigmoid function is applied to compute the probability of selecting the next item, following the approach suggested by recent literature (wu2023generic). The policy $\pi(a_{t}\mid s_{t})$ guiding action selection is defined as:
@@
[…]
@@
1 $\rightarrow v_{2}\rightarrow v_{3}\rightarrow v_{4}\rightarrow v_{5} \setminus$ 11
2 $v_{1}\rightarrow\_\rightarrow v_{3}\rightarrow v_{4}\rightarrow v_{5}$ $v_{1}$ 0
3 $v_{1}\rightarrow v_{2}\rightarrow\_\rightarrow v_{4}\rightarrow v_{5}$ $v_{1}$ 0
4 $v_{1}\rightarrow v_{2}\rightarrow v_{3}\rightarrow\_\rightarrow v_{5}$ $v_{1}$ 11
5 $v_{1}\rightarrow v_{2}\rightarrow v_{3}\rightarrow v_{4}\rightarrow$\_ $v_{1}\rightarrow v_{4}$ 11
Final - $v_{1}\rightarrow v_{4}\rightarrow v_{5}$ -
A Toy Example of the MDP process. To illustrate the execution process of our FCESR framework, consider a target session $v_{1}\rightarrow v_{2}\rightarrow v_{3}\rightarrow v_{4}\rightarrow v_{5}$”. The task involves traversing the session sequentially. Starting with an empty CF2 set $\mathcal{S}^{*}$, the agent decides at each step whether to include the current item in $\mathcal{S}^{*}$ based on the action $a_{t}$. At step 1, the agent begins with $v_{1}$ and chooses to include it in the explanation since $a_{1}=1$, updating $\mathcal{S}^{*}$ to $\{v_{1}\}$. At step 2, the agent moves to $v_{2}$ but decides not to include it since $a_{2}=0$, leaving $\mathcal{S}^{*}$ unchanged. In step 3, the agent processes $v_{3}$ and again decides not to include it in the explanation since $a_{3}=0$. At step 4, the agent evaluates $v_{4}$ and chooses to add it to $S^{*}$ since $a_{4}=1$, updating the explanation to $\{v_{1},v_{4}\}$. Finally, at step 5, the agent considers $v_{5}$ and includes it in the explanation since $a_{5}=1$, resulting in the final CF2 set $\mathcal{S}^{*}=\{v_{1},v_{4},v_{5}\}$. The action selection at each step is determined by a policy network, which integrates the embeddings of the current CF2. This example demonstrates how our framework effectively reduces the original session to its key components through a sequential decision-making process.